The sum of one-half, one-third and one-fourth of a number exceeds the number by 12. The number is
Found 2 solutions by ewatrrr, mananth:
Answer by ewatrrr(24785)


Hi
Let x represent the number
Write as we Read***
x/2 + x/3 + x/4 = x + 12
solving for x
6x +4x + 3x = 12x + 144 |multiplying each term by 12
x = 144
CHECKING our Answer
72 + 48 + 36 = 156 = 144 + 12

Answer by mananth(16946)

The sum of one-half, one-third and one-fourth of a number exceeds the number by 12. The number is
let the number be x
x/2 +x/3+x/4 = x+12
...
LCD = 12
(6x+4x+3x)/12=x+12
13x= 12(x+12)
13x=12x+144
13x-12x=144
x= 144
..